OpenStudy (anonymous):

17 years into minutes and show work please

OpenStudy (anonymous):

17 years into minutes and show work please

OpenStudy (anonymous):

1hr = 60 min 24hr = 24 * 60 = 1440 min (that is 1 day) 365 days = 365 * 1440 = 525600 min (that is 1 year) 17 years = 17 * 525600 = 8935200 min